A company has the data shown in the chart below concerning its forecast performance over the past five time periods.

Period

Actual Demand

Forecast Demand

Error

Absolute Value of Error

Absolute Percentage Error

1

345

320

25

2

320

10

3

335

350

4

340

-30

5

350

20



Calculate the running sum of forecast error (RSFE).
Answer #1

Period Error Absolute Error 320 25 10 Actual demand Forecast 345 320 335 310 370 -15 15 310 350 340 350 Percentage Error 7.24

Forecast of period 2 = 320 - 10 = 310

Actual demand for period 4 = 340 + (-30) = 310

Actual demand for period 5 = 350 + 20 = 370.

Do give a positive rating if you think this answer helped